OpenStudy (anonymous):

−4x + 10 = 2, which order of operations completely solves for x

OpenStudy (anonymous):

you mean whats x value, ,?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well , -4x + 10 = 2 - 4x + 10 - 10 = 2 - 10 - 4x = - 8 x = -8 / - 4 x = 2